# Semi-definite optimization of the measured relative entropies of quantum states and channels

### v2

@@ 6.2 Optimizing the measured relative entropy of channels
$D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})=\sup_{\Omega,\rho>0,\Theta\in\operatorname{Herm}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Theta\Gamma^{\mathcal{N}}]-\operatorname{Tr}[\Omega\Gamma^{\mathcal{M}}]+1:\\ \operatorname{Tr}[\rho]=1,\quad\operatorname{Tr}[H\rho]\leq E,\\ \Theta\leq P_{\ln}(\rho\otimes I,\Omega)\end{array}\right\},$ (88)
@@
Furthermore, when $\Gamma^{\mathcal{N}}$ and $\Gamma^{\mathcal{M}}$ are $d_{A}d_{B}\times d_{A}d_{B}$ matrices, the quantity $D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})$ can be efficiently calculated by means of a semi-definite program up to an error $\varepsilon$, by means of $O(\sqrt{\ln(1/\varepsilon)})$ linear matrix inequalities, each of size $2d_{A}d_{B}\times 2d_{A}d_{B}$.
@@
$=\sup_{\begin{subarray}{c}\Omega^{\prime}>0,\Theta^{\prime}\in\operatorname{Herm},\\ \rho_{R}\in\mathbb{D}(\mathcal{H}_{R}\mathcal{)},\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Theta^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}\rho_{R}^{1/2}]-\operatorname{Tr}[\Omega^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}\rho_{R}^{1/2}]+1:\\ \Theta^{\prime}\leq P_{\ln}(I,\Omega^{\prime})\end{array}\right\}$ (92)
$=\sup_{\begin{subarray}{c}\Omega^{\prime},\rho_{R}>0,\Theta^{\prime}\in\operatorname{Herm}\\ \operatorname{Tr}[\rho_{R}]=1\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\rho_{R}^{1/2}\Theta^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}]-\operatorname{Tr}[\rho_{R}^{1/2}\Omega^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}]+1:\\ \Theta^{\prime}\leq P_{\ln}(I,\Omega^{\prime})\end{array}\right\}$ (95)
$=\sup_{\begin{subarray}{c}\Omega,\rho_{R}>0,\Theta\in\operatorname{Herm}\\ \operatorname{Tr}[\rho_{R}]=1\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Theta\Gamma_{RB}^{\mathcal{N}}]-\operatorname{Tr}[\Omega\Gamma_{RB}^{\mathcal{M}}]+1:\\ \Theta\leq P_{\ln}(\rho\otimes I,\Omega)\end{array}\right\}.$ (98)
@@
$\Omega^{\prime}>0,\quad\Theta^{\prime}\in\operatorname{Herm}\quad\Leftrightarrow\quad\Omega>0,\quad\Theta\in\operatorname{Herm},$ (101)
@@
As such, we have rewritten $D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})$ in terms of the hypograph of $P_{\ln}$. By appealing to [FSP19, Theorem 2 and Theorem 3], it follows that this quantity can be efficiently calculated by means of semi-definite programming, with the complexity stated above.
@@ 7 Conclusion
Going forward from here, we note that further work in this direction could consider combining our findings here with those of [RSB24], the latter being about measured Rényi divergences under restricted forms of measurements. We also think it is a very interesting open question, related to those from [FS17, FS23], to determine semi-definite programs for various Rényi relative entropies of quantum channels, including those based on the sandwiched [MLDS+13, WWY14] and Petz–Rényi [Pet85, Pet86]  relative entropies. More generally, one could consider the same question for $\alpha$-$z$ Rényi relative entropies [AD15], and here we think the variational formulas from [MH24, Appendix B] could be useful in addressing this question.
@@ Appendix B $\alpha\to 1$ limit of the measured Rényi relative entropy
as noted in [RSB24, Lemma 2], and because

### v3

@@ 6.2 Optimizing the measured relative entropy of channels
$D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})=\sup_{\Omega,\rho>0,\Theta\in\mathbb{H}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Theta\Gamma^{\mathcal{N}}]-\operatorname{Tr}[\Omega\Gamma^{\mathcal{M}}]+1:\\ \operatorname{Tr}[\rho]=1,\quad\operatorname{Tr}[H\rho]\leq E,\\ \Theta\leq P_{\ln}(\rho\otimes I,\Omega)\end{array}\right\},$ (122)
@@
Furthermore, when $\Gamma^{\mathcal{N}}$ and $\Gamma^{\mathcal{M}}$ are $d_{A}d_{B}\times d_{A}d_{B}$ matrices, the quantity $D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})$ can be efficiently calculated by means of a semi-definite program up to an additive error $\varepsilon$, by means of $O(\sqrt{\ln(1/\varepsilon)})$ linear matrix inequalities, each of size $2d_{A}d_{B}\times 2d_{A}d_{B}$. Specifically, there exist $m,k\in\mathbb{N}$ such that $m+k=O(\sqrt{\ln(1/\varepsilon)})$ and the following inequality holds:
$\left|D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})-D_{H,E,m,k}^{M}(\mathcal{N}\|\mathcal{M})\right|\leq\varepsilon,$ (123)
where
$D_{H,E,m,k}^{M}(\mathcal{N}\|\mathcal{M})\coloneqq\\ \sup_{\begin{subarray}{c}\Omega,\rho>0,\Theta\in\mathbb{H},\\ T_{1},\ldots,T_{m}\in\mathbb{H},\\ Z_{0},\ldots,Z_{k}\in\mathbb{H}\end{subarray}}\left\{\begin{array}[]{c}\operatorname{Tr}\!\left[\Theta\Gamma^{\mathcal{N}}\right]-\operatorname{Tr}\!\left[\Omega\Gamma^{\mathcal{M}}\right]+1:\\ \operatorname{Tr}[\rho]=1,\>\operatorname{Tr}[H\rho]\leq E,\\ Z_{0}=\Omega,\,\sum_{j=1}^{m}w_{j}T_{j}=2^{-k}\Theta,\\ \left\{\begin{bmatrix}Z_{i}&Z_{i+1}\\ Z_{i+1}&\rho\otimes I\end{bmatrix}\geq 0\right\}_{i=0}^{k-1},\\ \left\{\begin{bmatrix}Z_{k}-\rho\otimes I-T_{j}&-\sqrt{t_{j}}T_{j}\\ -\sqrt{t_{j}}T_{j}&\rho\otimes I-t_{j}T_{j}\end{bmatrix}\geq 0\right\}_{j=1}^{m}\end{array}\right\}$ (124)
and, for all $j\in\left\{1,\ldots,m\right\}$, $w_{j}$ and $t_{j}$ are the weights and nodes, respectively, for the $m$-point Gauss–Legendre quadrature on the interval $\left[0,1\right]$.
@@
$=\sup_{\begin{subarray}{c}\Omega^{\prime}>0,\Theta^{\prime}\in\mathbb{H},\\ \rho_{R}\in\mathbb{D}(\mathcal{H}_{R}\mathcal{)},\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Theta^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}\rho_{R}^{1/2}]-\operatorname{Tr}[\Omega^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}\rho_{R}^{1/2}]+1:\\ \Theta^{\prime}\leq P_{\ln}(I,\Omega^{\prime})\end{array}\right\}$ (128)
$=\sup_{\begin{subarray}{c}\Omega^{\prime},\rho_{R}>0,\Theta^{\prime}\in\mathbb{H},\\ \operatorname{Tr}[\rho_{R}]=1,\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\rho_{R}^{1/2}\Theta^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{N}}]-\operatorname{Tr}[\rho_{R}^{1/2}\Omega^{\prime}\rho_{R}^{1/2}\Gamma_{RB}^{\mathcal{M}}]+1:\\ \Theta^{\prime}\leq P_{\ln}(I,\Omega^{\prime})\end{array}\right\}$ (131)
$=\sup_{\begin{subarray}{c}\Omega,\rho_{R}>0,\Theta\in\mathbb{H},\\ \operatorname{Tr}[\rho_{R}]=1,\\ \operatorname{Tr}[H_{A}\rho_{A}]\leq E\end{subarray}}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Theta\Gamma_{RB}^{\mathcal{N}}]-\operatorname{Tr}[\Omega\Gamma_{RB}^{\mathcal{M}}]+1:\\ \Theta\leq P_{\ln}(\rho\otimes I,\Omega)\end{array}\right\}.$ (134)
@@
$\Omega^{\prime}>0,\quad\Theta^{\prime}\in\mathbb{H}\quad\Leftrightarrow\quad\Omega>0,\quad\Theta\in\mathbb{H},$ (137)
@@
As such, we have rewritten $D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})$ in terms of the hypograph of $P_{\ln}$. By appealing to [FSP19, Theorem 2 and Theorem 3], it follows that this quantity can be efficiently calculated by means of semi-definite programming, with the complexity stated above. In more detail, following the proof of Proposition 7 closely and recalling the definition of $\operatorname{hyp}_{r_{m,k}}$ and its semi-definite representation in (66)–(70), observe that
$D_{H,E,m,k}^{M}(\mathcal{N}\|\mathcal{M})=\sup_{\Omega,\rho>0,\Theta\in\mathbb{H}}\left\{\begin{array}[]{c}\operatorname{Tr}\!\left[\Theta\Gamma^{\mathcal{N}}\right]-\operatorname{Tr}\!\left[\Omega\Gamma^{\mathcal{M}}\right]+1:\\ \operatorname{Tr}[\rho]=1,\>\operatorname{Tr}[H\rho]\leq E,\\ \Theta\leq P_{r_{m,k}}(\rho\otimes I,\Omega)\end{array}\right\}.$ (140)
By invoking the first part of Lemma 8, we conclude that
$\Theta\leq P_{\ln}(\rho\otimes I,\Omega)\quad\implies\quad\Theta-\varepsilon\left(\rho\otimes I\right)\leq P_{r_{m,k}}(\rho\otimes I,\Omega),$ (141)
so that
$D_{H,E}^{M}(\mathcal{N}\|\mathcal{M})$
$=\sup_{\Omega,\rho>0,\Theta\in\mathbb{H}}\left\{\begin{array}[]{c}\operatorname{Tr}\!\left[\Theta\Gamma^{\mathcal{N}}\right]-\operatorname{Tr}\!\left[\Omega\Gamma^{\mathcal{M}}\right]+1:\\ \operatorname{Tr}[\rho]=1,\>\operatorname{Tr}[H\rho]\leq E,\\ \Theta\leq P_{\ln}(\rho\otimes I,\Omega)\end{array}\right\}$ (145)
$\leq\sup_{\Omega,\rho>0,\Theta\in\mathbb{H}}\left\{\begin{array}[]{c}\operatorname{Tr}\!\left[\Theta\Gamma^{\mathcal{N}}\right]-\operatorname{Tr}\!\left[\Omega\Gamma^{\mathcal{M}}\right]+1:\\ \operatorname{Tr}[\rho]=1,\>\operatorname{Tr}[H\rho]\leq E,\\ \Theta-\varepsilon\left(\rho\otimes I\right)\leq P_{r_{m,k}}(\rho\otimes I,\Omega)\end{array}\right\}$ (149)
$=\sup_{\Omega,\rho>0,\Theta^{\prime}\in\mathbb{H}}\left\{\begin{array}[]{c}\operatorname{Tr}\!\left[\left(\Theta^{\prime}+\varepsilon\left(\rho\otimes I\right)\right)\Gamma^{\mathcal{N}}\right]-\operatorname{Tr}\!\left[\Omega\Gamma^{\mathcal{M}}\right]+1:\\ \operatorname{Tr}[\rho]=1,\>\operatorname{Tr}[H\rho]\leq E,\\ \Theta^{\prime}\leq P_{r_{m,k}}(\rho\otimes I,\Omega)\end{array}\right\}$ (153)
$=\sup_{\Omega,\rho>0,\Theta^{\prime}\in\mathbb{H}}\left\{\begin{array}[]{c}\operatorname{Tr}\!\left[\Theta^{\prime}\Gamma^{\mathcal{N}}\right]-\operatorname{Tr}\!\left[\Omega\Gamma^{\mathcal{M}}\right]+1:\\ \operatorname{Tr}[\rho]=1,\>\operatorname{Tr}[H\rho]\leq E,\\ \Theta^{\prime}\leq P_{r_{m,k}}(\rho\otimes I,\Omega)\end{array}\right\}+\varepsilon$ (157)
$=D_{H,E,m,k}^{M}(\mathcal{N}\|\mathcal{M})+\varepsilon.$ (158)
The second equality follows from the substitution $\Theta^{\prime}=\Theta-\varepsilon\left(\rho\otimes I\right)$ and the penultimate equality follows because
$\operatorname{Tr}\!\left[\left(\Theta^{\prime}+\varepsilon\left(\rho\otimes I\right)\right)\Gamma^{\mathcal{N}}\right]$ $=\operatorname{Tr}\!\left[\Theta^{\prime}\Gamma^{\mathcal{N}}\right]+\varepsilon\operatorname{Tr}\!\left[\left(\rho\otimes I\right)\Gamma^{\mathcal{N}}\right]$ (159)
$=\operatorname{Tr}\!\left[\Theta^{\prime}\Gamma^{\mathcal{N}}\right]+\varepsilon,$ (160)
given that $\operatorname{Tr}\!\left[\left(\rho\otimes I\right)\Gamma^{\mathcal{N}}\right]=\operatorname{Tr}\!\left[\mathcal{N}(\rho^{T})\right]=1$, where the second equality follows from [KW20, Eq. (4.2.12)]. Now by invoking the second part of Lemma 8, we conclude that
$\Theta\leq P_{r_{m,k}}(\rho\otimes I,\Omega)\quad\implies\quad\Theta-\varepsilon\left(\rho\otimes I\right)\leq P_{\ln}(\rho\otimes I,\Omega),$ (161)
so that
$D_{H,E,m,k}^{M}(\mathcal{N}\|\mathcal{M})$
$=\sup_{\Omega,\rho>0,\Theta\in\mathbb{H}}\left\{\begin{array}[]{c}\operatorname{Tr}\!\left[\Theta\Gamma^{\mathcal{N}}\right]-\operatorname{Tr}\!\left[\Omega\Gamma^{\mathcal{M}}\right]+1:\\ \operatorname{Tr}[\rho]=1,\>\operatorname{Tr}[H\rho]\leq E,\\ \Theta\leq P_{r_{m,k}}(\rho\otimes I,\Omega)\end{array}\right\}$ (165)
$\leq\sup_{\Omega,\rho>0,\Theta\in\mathbb{H}}\left\{\begin{array}[]{c}\operatorname{Tr}\!\left[\Theta\Gamma^{\mathcal{N}}\right]-\operatorname{Tr}\!\left[\Omega\Gamma^{\mathcal{M}}\right]+1:\\ \operatorname{Tr}[\rho]=1,\>\operatorname{Tr}[H\rho]\leq E,\\ \Theta-\varepsilon\left(\rho\otimes I\right)\leq P_{\ln}(\rho\otimes I,\Omega)\end{array}\right\}$ (169)
$=\sup_{\Omega,\rho>0,\Theta^{\prime}\in\mathbb{H}}\left\{\begin{array}[]{c}\operatorname{Tr}\!\left[\left(\Theta^{\prime}+\varepsilon\left(\rho\otimes I\right)\right)\Gamma^{\mathcal{N}}\right]-\operatorname{Tr}\!\left[\Omega\Gamma^{\mathcal{M}}\right]+1:\\ \operatorname{Tr}[\rho]=1,\>\operatorname{Tr}[H\rho]\leq E,\\ \Theta^{\prime}\leq P_{\ln}(\rho\otimes I,\Omega)\end{array}\right\}$ (173)
$=\sup_{\Omega,\rho>0,\Theta^{\prime}\in\mathbb{H}}\left\{\begin{array}[]{c}\operatorname{Tr}\!\left[\Theta^{\prime}\Gamma^{\mathcal{N}}\right]-\operatorname{Tr}\!\left[\Omega\Gamma^{\mathcal{M}}\right]+1:\\ \operatorname{Tr}[\rho]=1,\>\operatorname{Tr}[H\rho]\leq E,\\ \Theta^{\prime}\leq P_{\ln}(\rho\otimes I,\Omega)\end{array}\right\}+\varepsilon$ (177)
$=D_{H,E,m,k}^{M}(\mathcal{N}\|\mathcal{M})+\varepsilon.$ (178)
Combining (158) and (178), we conclude (123).
@@ 7 Conclusion
Going forward from here, we note that further work in this direction could consider combining our findings here with those of [RSB25], the latter being about measured Rényi divergences under restricted forms of measurements. We also think it is a very interesting open question, related to those from [FS17, FS23], to determine semi-definite programs for various Rényi relative entropies of quantum channels, including those based on the sandwiched [MLDS+13, WWY14] and Petz–Rényi [Pet85, Pet86] relative entropies. More generally, one could consider the same question for $\alpha$-$z$ Rényi relative entropies [AD15], and here we think the variational formulas from [MH24, Appendix B] could be useful in addressing this question.
@@ Appendix B $\alpha\to 1$ limit of the measured Rényi relative entropy
as noted in [RSB25, Lemma 2], and because